OpenStudy (anonymous):

Line BC contains points B (4, −5) and C (3, 2). Line DE contains points D (2, 0) and E (9, 1). Lines BC and DE are parallel perpendicular neither

OpenStudy (triciaal):

 this is an approximation definitely not parallel perpendicular lines the product of the slopes =-1 slope = change in y/change in x slope of bc 2-(-5)/3-4 = -7 slope of de 1-0/9-2 = 1/7 product of slopes -7*1/7 =-1 lines are perpendicular